OpenStudy (anonymous):

Find the equation of the line that is parallel to the line formed by the equation 4x - 3y = -9 and passes through the point (-18, 1)

OpenStudy (kkutie7):

Solve for y. Slope of the new line should be the same as this one. Plug into this y=slope(x-x1)+y1

OpenStudy (whpalmer4):

Rearrange your original line's equation by solving it for \(y\). You'll end up with the equation in slope-intercept form: \[y = mx+b\]where \(m\) is the slope and \(b\) is the value of the y-intercept. A parallel line will have an equal slope. Use the point-slope formula \[y-y_1 = m(x-x_1)\]to write the equation of a line with slope \(m\) (as found above) that passes through the point \((x_1, y_1)\) (given by the problem statement). Rearrange the resulting equation to suit your needs, if necessary.